OpenStudy (anonymous):

A clock takes 12 seconds to strike 12 times at 12 o'clock. how long does it take to strike 6 times at 6 o'clock?

OpenStudy (aum):

We start timing when we hear the first strike. When the clock strikes 12 times, we have timed 11 strikes after the first strike. That takes 12 seconds. When the clock strikes 6 times, we have to time 5 strikes after the first strike. From proportions we can say it will take 5 * 12 / 11 = 60/11 = 5 5/11 seconds.

OpenStudy (aum):

12 seconds / 11 strikes = x seconds / 5 strikes x = 5 * 12 / 11 = 60/11 = 5 5/11 seconds.